OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. (sin x/1-cos x) + (sin x/ 1+cos x)= 2csc x

OpenStudy (anonymous):

@SithsAndGiggles @surjithayer

OpenStudy (anonymous):

i think i figured it out

OpenStudy (anonymous):

sin x/(1-cosx)=cot(x/2) sin x/(1+cosx)=tan(x/2) cot(x/2)+tan(x/2)=2 cscx is this how to do it?

OpenStudy (anonymous):

\[\frac{ \sin x }{ 1-\cos x }+\frac{ \sin x }{ 1+\cos x }=\sin x \left( \frac{ 1 }{ 1-\cos x } +\frac{ 1 }{ 1+\cos x }\right)\] \[=\sin x \left( \frac{ 1+\cos x+1-\cos x }{ 1-\cos ^2x } \right)=\frac{ 2 \sin x }{ \sin ^2x }=\frac{ 2 }{ \sin x }=?\]

OpenStudy (anonymous):

2/sinx= 2csc x

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

can you help me with one more? @surjithayer

OpenStudy (anonymous):

i will try.

OpenStudy (anonymous):

-tan^2x+sec^2x=1

OpenStudy (anonymous):

we know \[\sin ^2x+\cos ^2x=1\] divide both sides by \[\cos ^2x\]

OpenStudy (anonymous):

i think now you can solve.

OpenStudy (anonymous):

sin^2x=1/cos^2x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ \sin ^2x }{ \cos ^2x }+\frac{ \cos ^2x }{ \cos ^2x }=\frac{ 1 }{ \cos ^2x }\] \[\tan ^2x+1=\sec ^2x, \]

OpenStudy (anonymous):

how does that equal to one? @surjithayer

OpenStudy (anonymous):

\[1=\sec ^2x-\tan ^2x\]

OpenStudy (anonymous):

or \[-\tan ^2x+\sec ^2x=1\]

OpenStudy (anonymous):

ooooh I see now!! sorry :p thank you!

OpenStudy (anonymous):

yw